# Talk2Move: Reinforcement Learning for Text-Instructed Object-Level Geometric Transformation in Scenes

### v1

@@ Abstract
The work was done during an internship at AWS Agentic AI.
@@ 1 Introduction
Current graphics-based editing methods [dragyourgan, mou2023dragondiffusion, 3ditscene, li2024omnidrag] often rely on 2D/3D primitives as control signals to assist scene editing. This requires human intervention and domain expertise in 2D/3D generation, making it not user-friendly nor in natural form of interaction for general users. Moreover, these approaches typically demands large amounts of high-quality paired data for SFT training, yet such data are rare and expensive to obtain. Spatial-manipulation pairs are particularly scarce—video datasets or 3D simulations provide only limited examples and are costly to scale. Top-down modeling of image content exists [gallagher2015happened, reed2015deep] which requires small training data but their performances are not meeting the practical requirement. Moreover, SFT-based methods rely on pixel-level MSE loss that struggles to disentangle objects from scenes, leading to limited spatial control.
@@ 2 Related Work
Drag-based Spatial Manipulation. Drag-based editing enables explicit object control through handle–target interactions. 2D methods [dragyourgan, mou2023dragondiffusion, dragyournoise, shi2024dragdiffusion, shin2024instantdrag, lu2024regiondrag] rely on point tracking or feature-space guidance to iteratively move objects, but they require manual specification of control points and struggle with complex, high-level instructions.
@@
To introduce geometric priors, 3D-aware approaches [3dfixup, 3ditscene, imagesculpting] lift objects into 3D space and perform transformations on reconstructed 3D representation. However, these pipelines involve multi-stage 2D-to-3D lifting and rendering, leading to high complexity, cumulative errors, and reduced image fidelity. Overall, both 2D and 3D dragging methods depend on explicit spatial primitives and user intervention, making them less suitable for natural scene-level object manipulation.
@@
Text-based Image Editing. Recent image generation and editing methods fall broadly into two families: diffusion-based models and LLM/VLM-connected frameworks. Diffusion-based methods [stablediffusion, ramesh2021zero] generate images via denoising, with recent flow-matching models such as Flux and Flux.1 Kontext [fluxkontext, blackforestlabs2023flux] improving convergence and edit consistency through in-context learning. However, relying solely on text tokenizers limits their ability to follow fine-grained spatial instructions.
@@
Another line of work integrates pre-trained LLMs/VLMs with diffusion decoders via lightweight adapters. Methods such as MetaQueries [metaqueries], Seed-X [seedx], and Emu2 [emu2], as well as unified transformer frameworks like MoT and Bagel [mot, bagel], enhance prompt following but often struggle to preserve visual coherence when only ViT-based encoders are used. QwenImageEdit [qwenimage] addresses this by combining VLM semantic features with VAE reconstructive features for stronger alignment between instruction following and visual fidelity. We adopt QwenImageEdit as our backbone for spatial manipulation tasks.
@@
Reinforcement Learning for Visual Generation. Reinforcement learning for diffusion-based generation typically formulates the denoising process as a Markov Decision Process (MDP). Reward-Weighted Regression [rwr] models generation as a one-step MDP and shifts the image distribution toward high-reward samples. DDPO [ddpo] extends this to a multi-step MDP by treating each denoising step as an action, making it well-suited for deterministic samplers without requiring repeated responses from the same query. Recent works adapt GRPO to diffusion models. FlowGRPO [flowgrpo] and DanceGRPO [dancegrpo] introduce stochastic drift into DDPM and Rectified Flow samplers, enabling multiple rollout paths around the deterministic trajectory. Subsequent improvements [mixgrpo, branchgrpo, g2rpo] reduce training cost via sliding-window optimization or tree-based trajectory reorganization with reward backpropagation. Although effective, these methods remain computationally expensive during sampling. In this work, we further improve time efficiency by introducing an early-exit mechanism to accelerate rollout generation. TempFlow-GRPO [he2025tempflow] shows that temporal reweighting improves general GRPO stability. Our work instead focuses on identifying task-dependent transformation stages for object-level geometry. RL-based image editing methods [li2024instructrl4pix, editscore, li2025uniworld] exist but they are not focused on the object transformation tasks.
@@ 4 Geometric Transformation Data Curation
Reference Image Generation. We prompt a large language model (LLM) to generate textual scene descriptions that specify realistic scenes with a background and multiple sparsely distributed movable objects, and then employ open-source text-to-image models [fluxkontext, qwenimage] to synthesize the corresponding reference images.
@@
Target Image Synthesis for SFT Cold Start. Current image editing backbones struggle with precise spatial manipulation, so we first apply an SFT cold-start to learn basic spatial-editing priors from a small amount of high-quality ground-truth data. However, constructing paired spatial-editing data is challenging, as real-world datasets rarely provide matched before–after examples for specific geometric transformations. To address this, we synthesize supervision with tailored strategies for translation, rotation, and resizing. For object translation and rotation, we use the API-based video generation model [gao2025seedream] to simulate physically plausible object motion in diverse scenes. Conditioned on a reference image as the first frame and a manipulation prompt, the model generates short videos depicting the desired transformation. After filtering by spatial accuracy and visual consistency, we obtain 800 valid translation pairs and 43 high-quality rotation pairs. For object resizing, where current video generation models perform poorly, we instead rely on open-source image editing models [qwenimage, fluxkontext] to synthesize coarse up/down scaling. The edited images are paired with their references and filtered via perceptual matching, yielding 110 diverse resizing pairs.
@@ 5.1 Preliminary
Flow-based GRPO [flowgrpo, dancegrpo] extends GRPO to flow-based generation by modeling the denoising trajectory as a Markov Decision Process (MDP). During inference, the diffusion model usually starts from a random Gaussian noise $\mathbf{x}_{T}$ and performs a reverse denoising process over $T$ steps. At each step, the model predicts the noise component conditioned on the input signals (e.g., text or image features) and progressively removes it, gradually reconstructing a clean image $\mathbf{x}_{0}$. Each denoising step can be viewed as a Reinforcement Learning (RL) state $s_{t}=(\mathbf{c},\mathbf{x}_{t})$, where $\mathbf{x}_{t}$ is the noisy latent and $\mathbf{c}$ is the conditioning input. The action corresponds to predicting the previous latent $a_{t}=\mathbf{x}_{t-1}$, and the policy is defined as the conditional generation step: $\pi(a_{t}|s_{t})=p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c}).$ The entire $T$-step denoising trajectory thus forms an MDP episode. At each step, the log probability represents the likelihood of sampling the previous latent $\mathbf{x}_{t-1}$ under the model-predicted conditional Gaussian transition $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})$.
@@ 5.2 Training Pipeline
RL implementation. To achieve data-efficient and spatially accurate scene manipulation, Talk2Move adopts a GRPO-style reinforcement learning framework tailored for diffusion policies. Starting from an initial noise sample, we follow the flow-based GRPO paradigm [flowgrpo, dancegrpo], injecting stochastic perturbations at each diffusion step to form groups of diverse sampling trajectories. These rollouts capture a rich set of spatial transformations conditioned on the input prompt and diffusion dynamics. Spatially grounded rewards are provided by off-the-shelf specialist models that disentangles objects explicitly from the scene and evaluates object-level transformations, from which we compute group-relative advantages to update model parameters via policy gradients.
@@ 5.3 Efficient GRPO with Early Exit
Off-policy Step Evaluation. Inspired by reasoning LLMs [wang2025beyond], which emphasize tokens with high intrinsic uncertainty, we propose to focus on diffusion steps exhibiting high extrinsic uncertainty. Since the noise variance in diffusion is fixed by the scheduler, we measure uncertainty through the variance of rollout rewards at each step. High reward variance indicates strong exploration potential and thus higher learning utility, analogous to high-entropy tokens in language RL. In addition, according to MixGRPO [mixgrpo] and our empirical findings, early denoising steps largely determine global layout, while later steps refine fine-grained details. Therefore, the step importance is defined as the reward variance of rollouts when exiting from each denoising step. In practice, starting from step 0, we incrementally perturb small sets of denoising steps on a few input images (2–4) and compute rollout rewards. This efficient calibration, requiring only a single GPU, yields a task-specific, step-aware reward distribution. The optimal exit step is then selected as the last perturbed step achieving the maximum reward variance.
@@
Flow-based GRPO approaches apply ODE updates at every denoising step, resulting in a time complexity of $T(t_{\text{sample}}+t_{\text{optim}})$ for $T$ steps. The sliding-window mechanism in [mixgrpo] alleviates this by optimizing only a subset of steps per iteration. We further enhance efficiency by directly denoising from the exit step $K$ to the final step $T$ using model predictions (see Fig.˜3), effectively shortening the trajectory. This reduces total computation to approximately $K(t_{\text{sample}}+t_{\text{optim}})$, where $K\leq T$. As long as $K<T$, our method achieves consistently higher training efficiency while maintaining comparable task performance.
Figure 3: Three types of step sampling: (a) is the full sampling and optimizing GRPO [flowgrpo, dancegrpo]; subsequent methods [mixgrpo, branchgrpo] as in (b), use a sliding window (yellow) to reduce the optimizing steps per iteration; (c) our work introduces step-wise active sampling that select the informative steps (red) and use shortcuts to bypass the rest of the steps, reducing both the sampling and optimizing time.
@@ 5.4 Spatially Grounded Reward Design
The choice of reward model is one of the most important factors in GRPO training, as it provides critical information for the policy update direction. Previous works [flowgrpo, editscore] employ holistic, image-level metrics such as aesthetic rewards, CLIP-based alignment, or VLMs fine-tuned for common editing tasks. These rewards serve as a coarse measure of overall image quality or human preference. We extend the reward model to fine-grained, spatially grounded rewards tailored for spatial manipulation editing.
To evaluate the correctness of geometric object transformation, we first localize the prompted object in both the reference and edited images using text-driven segmentation [sam, groundingdino] to obtain segmentation masks and bounding boxes. Each manipulation task defines a specific spatial reward. For object translation, we compute the relative displacement between the reference and edited object centers following the GenEval [geneval] protocol. We also employ depth estimation model [depthanything, wang2025moge] to obtain object depth value before and after editing to reward forward and backward movements. For rotation, we use Orient-Anything [orientanything] to estimate orientations for object crops from the scene image and evaluate alignment along the prompted axis, direction, and angle. For resizing, we compare the normalized size ratio differences of reference and edited object bounding boxes. Although each reward differs in formulation, all share a unified goal, which is to measure the alignment between the prompted and achieved spatial transformations in normalized coordinate space. These task-specific rewards provide fine-grained feedback to guide policy updates toward spatially accurate editing behavior.
@@ 6 Experiments
GPT-Image-1 [openai2024gptimage1] 0.5416 64.29% 26.25% 0.4293 2.33% 6.25% 0.3501 5.08% 1.39%
Flux-Kontext [fluxkontext] 0.0499 4.41% 1.25% 0.4259 6.82% 16.25% 0.4318 1.67% 15.28%
Bagel [bagel] 0.1705 14.49% 2.50% 0.3240 13.64% 1.25% 0.4790 0.00% 8.33%
QwenImageEdit [qwenimage] 0.2551 32.86% 12.50% 0.4129 9.30% 7.50% 0.4203 7.50% 11.11%
@@
Table 2: Quantitative comparison of object transformation tasks on curated real images from OpenImages-V6 [OpenImages2] in terms of editing accuracy, average translation distance and editing errors.
@@
GPT-Image-1 [openai2024gptimage1] 0.3237 23.08% 0.1774 25.00% 0.6898 7.14%
Flux-Kontext [fluxkontext] 0.1992 26.92% 0.2567 6.67% 0.7120 1.79%
Bagel [bagel] 0.2258 15.38% 0.3700 6.25% 0.7159 3.57%
QwenImageEdit [qwenimage] 0.3725 42.31% 0.2101 25.00% 0.6048 1.79%
@@ 6.1 Implementation Details
For cold start, we train Qwen-Image-Edit [qwenimage] with rank-64 LoRA layers for 3,000 iterations and with a learning rate of 1e-4. For GRPO training, we build upon the FlowGRPO [flowgrpo] baseline, with sample noise level of 1.0 and a clip range of 2e-4, on 16 H200 GPU server. The total training time for each subtask is $\sim 160$ GPU hours.
@@ 6.2 Evaluation Metrics
Translation. We report the Translation Distance (Trans. Dist.) that measures the relative distance between object center points in the reference and edited image; and the translation Accuracy (Acc.) that measures the success rate. An edit is considered successful only if it passes all four criteria: 1) correct movement: following GenEval [geneval], we verify whether the detected object in the edited image has moved in the specified direction relative to its position in the reference image; 2) object identity preservation: the edited object must remain the same object (CLIP similarity > 0.75); 3) no duplication: the object should no longer appear at its original location (similarity < 0.95); 4) scene preservation: the overall scene should remain consistent (background L1 distance $\leq 0.2$).
@@ 6.3 Quantitative Evaluation
We show the quantitative comparison with state-of-the-art image editing methods, including diffusion-based editing [fluxkontext], unified model [bagel], our baseline [qwenimage] and GPT-Image-1. Tab.˜1 reports Talk2Move’s result on our synthetic evaluation benchmark for three geometric transformation tasks, featuring 100 input samples under each transformation type. Our model outperforms open-sourced baselines among most metrics. We also sample 85 real images from OpenImagesV6 [OpenImages2] for evaluation in Tab.˜2 and results prove our model’s effectiveness on real images.

### v2

@@ Abstract
The work was done during an internship at AWS Agentic AI. Jing Tan is now with the Chinese University of Hong Kong. Correspondence to ozhaozha@amazon.com
@@ 1 Introduction
Current graphics-based editing methods [28, 25, 43, 16] often rely on 2D/3D primitives as control signals to assist scene editing. This requires human intervention and domain expertise in 2D/3D generation, making it not user-friendly nor in natural form of interaction for general users. Moreover, these approaches typically demands large amounts of high-quality paired data for SFT training, yet such data are rare and expensive to obtain. Spatial-manipulation pairs are particularly scarce—video datasets or 3D simulations provide only limited examples and are costly to scale. Top-down modeling of image content exists [5, 31] which requires small training data but their performances are not meeting the practical requirement. Moreover, SFT-based methods rely on pixel-level MSE loss that struggles to disentangle objects from scenes, leading to limited spatial control.
@@ 2 Related Work
Drag-based Spatial Manipulation. Drag-based editing enables explicit object control through handle–target interactions. 2D methods [28, 25, 20, 33, 34, 23] rely on point tracking or feature-space guidance to iteratively move objects, but they require manual specification of control points and struggle with complex, high-level instructions.
@@
To introduce geometric priors, 3D-aware approaches [3, 43, 42] lift objects into 3D space and perform transformations on reconstructed 3D representation. However, these pipelines involve multi-stage 2D-to-3D lifting and rendering, leading to high complexity, cumulative errors, and reduced image fidelity. Overall, both 2D and 3D dragging methods depend on explicit spatial primitives and user intervention, making them less suitable for natural scene-level object manipulation.
@@
Text-based Image Editing. Recent image generation and editing methods fall broadly into two families: diffusion-based models and LLM/VLM-connected frameworks. Diffusion-based methods [32, 30] generate images via denoising, with recent flow-matching models such as Flux and Flux.1 Kontext [12, 13] improving convergence and edit consistency through in-context learning. However, relying solely on text tokenizers limits their ability to follow fine-grained spatial instructions.
@@
Another line of work integrates pre-trained LLMs/VLMs with diffusion decoders via lightweight adapters. Methods such as MetaQueries [27], Seed-X [7], and Emu2 [35], as well as unified transformer frameworks like MoT and Bagel [19, 4], enhance prompt following but often struggle to preserve visual coherence when only ViT-based encoders are used. QwenImageEdit [39] addresses this by combining VLM semantic features with VAE reconstructive features for stronger alignment between instruction following and visual fidelity. We adopt QwenImageEdit as our backbone for spatial manipulation tasks.
@@
Reinforcement Learning for Visual Generation. Reinforcement learning for diffusion-based generation typically formulates the denoising process as a Markov Decision Process (MDP). Reward-Weighted Regression [29] models generation as a one-step MDP and shifts the image distribution toward high-reward samples. DDPO [2] extends this to a multi-step MDP by treating each denoising step as an action, making it well-suited for deterministic samplers without requiring repeated responses from the same query. Recent works adapt GRPO to diffusion models. FlowGRPO [21] and DanceGRPO [40] introduce stochastic drift into DDPM and Rectified Flow samplers, enabling multiple rollout paths around the deterministic trajectory. Subsequent improvements [14, 17, 44] reduce training cost via sliding-window optimization or tree-based trajectory reorganization with reward backpropagation. Although effective, these methods remain computationally expensive during sampling. In this work, we further improve time efficiency by introducing an early-exit mechanism to accelerate rollout generation. TempFlow-GRPO [9] shows that temporal reweighting improves general GRPO stability. Our work instead focuses on identifying task-dependent transformation stages for object-level geometry. RL-based image editing methods [15, 24, 18] exist but they are not focused on the object transformation tasks.
@@ 4 Geometric Transformation Data Curation
Reference Image Generation. We prompt a large language model (LLM) to generate textual scene descriptions that specify realistic scenes with a background and multiple sparsely distributed movable objects, and then employ open-source text-to-image models [12, 39] to synthesize the corresponding reference images.
@@
Target Image Synthesis for SFT Cold Start. Current image editing backbones struggle with precise spatial manipulation, so we first apply an SFT cold-start to learn basic spatial-editing priors from a small amount of high-quality ground-truth data. However, constructing paired spatial-editing data is challenging, as real-world datasets rarely provide matched before–after examples for specific geometric transformations. To address this, we synthesize supervision with tailored strategies for translation, rotation, and resizing. For object translation and rotation, we use the API-based video generation model [6] to simulate physically plausible object motion in diverse scenes. Conditioned on a reference image as the first frame and a manipulation prompt, the model generates short videos depicting the desired transformation. After filtering by spatial accuracy and visual consistency, we obtain 800 valid translation pairs and 43 high-quality rotation pairs. For object resizing, where current video generation models perform poorly, we instead rely on open-source image editing models [39, 12] to synthesize coarse up/down scaling. The edited images are paired with their references and filtered via perceptual matching, yielding 110 diverse resizing pairs.
@@ 5.1 Preliminary
Flow-based GRPO [21, 40] extends GRPO to flow-based generation by modeling the denoising trajectory as a Markov Decision Process (MDP). During inference, the diffusion model usually starts from a random Gaussian noise $\mathbf{x}_{T}$ and performs a reverse denoising process over $T$ steps. At each step, the model predicts the noise component conditioned on the input signals (e.g., text or image features) and progressively removes it, gradually reconstructing a clean image $\mathbf{x}_{0}$. Each denoising step can be viewed as a Reinforcement Learning (RL) state $s_{t}=(\mathbf{c},\mathbf{x}_{t})$, where $\mathbf{x}_{t}$ is the noisy latent and $\mathbf{c}$ is the conditioning input. The action corresponds to predicting the previous latent $a_{t}=\mathbf{x}_{t-1}$, and the policy is defined as the conditional generation step: $\pi(a_{t}|s_{t})=p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c}).$ The entire $T$-step denoising trajectory thus forms an MDP episode. At each step, the log probability represents the likelihood of sampling the previous latent $\mathbf{x}_{t-1}$ under the model-predicted conditional Gaussian transition $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})$.
@@ 5.2 Training Pipeline
RL implementation. To achieve data-efficient and spatially accurate scene manipulation, Talk2Move adopts a GRPO-style reinforcement learning framework tailored for diffusion policies. Starting from an initial noise sample, we follow the flow-based GRPO paradigm [21, 40], injecting stochastic perturbations at each diffusion step to form groups of diverse sampling trajectories. These rollouts capture a rich set of spatial transformations conditioned on the input prompt and diffusion dynamics. Spatially grounded rewards are provided by off-the-shelf specialist models that disentangles objects explicitly from the scene and evaluates object-level transformations, from which we compute group-relative advantages to update model parameters via policy gradients.
@@ 5.3 Efficient GRPO with Early Exit
Off-policy Step Evaluation. Inspired by reasoning LLMs [37], which emphasize tokens with high intrinsic uncertainty, we propose to focus on diffusion steps exhibiting high extrinsic uncertainty. Since the noise variance in diffusion is fixed by the scheduler, we measure uncertainty through the variance of rollout rewards at each step. High reward variance indicates strong exploration potential and thus higher learning utility, analogous to high-entropy tokens in language RL. In addition, according to MixGRPO [14] and our empirical findings, early denoising steps largely determine global layout, while later steps refine fine-grained details. Therefore, the step importance is defined as the reward variance of rollouts when exiting from each denoising step. In practice, starting from step 0, we incrementally perturb small sets of denoising steps on a few input images (2–4) and compute rollout rewards. This efficient calibration, requiring only a single GPU, yields a task-specific, step-aware reward distribution. The optimal exit step is then selected as the last perturbed step achieving the maximum reward variance.
@@
Flow-based GRPO approaches apply ODE updates at every denoising step, resulting in a time complexity of $T(t_{\text{sample}}+t_{\text{optim}})$ for $T$ steps. The sliding-window mechanism in [14] alleviates this by optimizing only a subset of steps per iteration. We further enhance efficiency by directly denoising from the exit step $K$ to the final step $T$ using model predictions (see Fig.˜3), effectively shortening the trajectory. This reduces total computation to approximately $K(t_{\text{sample}}+t_{\text{optim}})$, where $K\leq T$. As long as $K<T$, our method achieves consistently higher training efficiency while maintaining comparable task performance.
Figure 3: Three types of step sampling: (a) is the full sampling and optimizing GRPO [21, 40]; subsequent methods [14, 17] as in (b), use a sliding window (yellow) to reduce the optimizing steps per iteration; (c) our work introduces step-wise active sampling that select the informative steps (red) and use shortcuts to bypass the rest of the steps, reducing both the sampling and optimizing time.
@@ 5.4 Spatially Grounded Reward Design
The choice of reward model is one of the most important factors in GRPO training, as it provides critical information for the policy update direction. Previous works [21, 24] employ holistic, image-level metrics such as aesthetic rewards, CLIP-based alignment, or VLMs fine-tuned for common editing tasks. These rewards serve as a coarse measure of overall image quality or human preference. We extend the reward model to fine-grained, spatially grounded rewards tailored for spatial manipulation editing.
To evaluate the correctness of geometric object transformation, we first localize the prompted object in both the reference and edited images using text-driven segmentation [10, 22] to obtain segmentation masks and bounding boxes. Each manipulation task defines a specific spatial reward. For object translation, we compute the relative displacement between the reference and edited object centers following the GenEval [8] protocol. We also employ depth estimation model [41, 36] to obtain object depth value before and after editing to reward forward and backward movements. For rotation, we use Orient-Anything [38] to estimate orientations for object crops from the scene image and evaluate alignment along the prompted axis, direction, and angle. For resizing, we compare the normalized size ratio differences of reference and edited object bounding boxes. Although each reward differs in formulation, all share a unified goal, which is to measure the alignment between the prompted and achieved spatial transformations in normalized coordinate space. These task-specific rewards provide fine-grained feedback to guide policy updates toward spatially accurate editing behavior.
@@ 6 Experiments
GPT-Image-1 [26] 0.5416 64.29% 26.25% 0.4293 2.33% 6.25% 0.3501 5.08% 1.39%
Flux-Kontext [12] 0.0499 4.41% 1.25% 0.4259 6.82% 16.25% 0.4318 1.67% 15.28%
Bagel [4] 0.1705 14.49% 2.50% 0.3240 13.64% 1.25% 0.4790 0.00% 8.33%
QwenImageEdit [39] 0.2551 32.86% 12.50% 0.4129 9.30% 7.50% 0.4203 7.50% 11.11%
@@
Table 2: Quantitative comparison of object transformation tasks on curated real images from OpenImages-V6 [11] in terms of editing accuracy, average translation distance and editing errors.
@@
GPT-Image-1 [26] 0.3237 23.08% 0.1774 25.00% 0.6898 7.14%
Flux-Kontext [12] 0.1992 26.92% 0.2567 6.67% 0.7120 1.79%
Bagel [4] 0.2258 15.38% 0.3700 6.25% 0.7159 3.57%
QwenImageEdit [39] 0.3725 42.31% 0.2101 25.00% 0.6048 1.79%
@@ 6.1 Implementation Details
For cold start, we train Qwen-Image-Edit [39] with rank-64 LoRA layers for 3,000 iterations and with a learning rate of 1e-4. For GRPO training, we build upon the FlowGRPO [21] baseline, with sample noise level of 1.0 and a clip range of 2e-4, on 16 H200 GPU server. The total training time for each subtask is $\sim 160$ GPU hours.
@@ 6.2 Evaluation Metrics
Translation. We report the Translation Distance (Trans. Dist.) that measures the relative distance between object center points in the reference and edited image; and the translation Accuracy (Acc.) that measures the success rate. An edit is considered successful only if it passes all four criteria: 1) correct movement: following GenEval [8], we verify whether the detected object in the edited image has moved in the specified direction relative to its position in the reference image; 2) object identity preservation: the edited object must remain the same object (CLIP similarity > 0.75); 3) no duplication: the object should no longer appear at its original location (similarity < 0.95); 4) scene preservation: the overall scene should remain consistent (background L1 distance $\leq 0.2$).
@@ 6.3 Quantitative Evaluation
We show the quantitative comparison with state-of-the-art image editing methods, including diffusion-based editing [12], unified model [4], our baseline [39] and GPT-Image-1. Tab.˜1 reports Talk2Move’s result on our synthetic evaluation benchmark for three geometric transformation tasks, featuring 100 input samples under each transformation type. Our model outperforms open-sourced baselines among most metrics. We also sample 85 real images from OpenImagesV6 [11] for evaluation in Tab.˜2 and results prove our model’s effectiveness on real images.